# Latent Topic Synthesis: Leveraging LLMs for Electoral Ad Analysis

### v1

@@ Abstract
Social media platforms play a pivotal role in shaping political discourse, but analyzing their vast and rapidly evolving content remains a major challenge. We introduce an end-to-end framework for automatically generating an interpretable topic taxonomy from an unlabeled corpus. By combining unsupervised clustering with prompt-based labeling, our method leverages large language models (LLMs) to iteratively construct a taxonomy without requiring seed sets or domain expertise. We apply this framework to a large corpus of Meta (previously known as Facebook) political ads from the month ahead of the 2024 U.S. Presidential election. Our approach uncovers latent discourse structures, synthesizes semantically rich topic labels, and annotates topics with moral framing dimensions. We show quantitative and qualitative analyses to demonstrate the effectiveness of our framework. Our findings reveal that voting and immigration ads dominate overall spending and impressions, while abortion and election-integrity achieve disproportionate reach. Funding patterns are equally polarized: economic appeals are driven mainly by conservative PACs, abortion messaging splits between pro- and anti-rights coalitions, and crime-and-justice campaigns are fragmented across local committees. The framing of these appeals also diverges—abortion ads emphasize liberty/oppression rhetoric, while economic messaging blends care/harm, fairness/cheating, and liberty/oppression narratives. Topic salience further reveals strong correlations between moral foundations and issues, such as fairness/cheating with crime/justice and loyalty/betrayal with immigration. Demographic targeting also emerges: younger Floridians are shown affordable-housing ads while older groups receive abortion-focused appeals; Montana males are disproportionately targeted with environmental and liberty-oriented messaging, whereas Virginia males are mobilized around crime/justice and voting. This work supports scalable, interpretable analysis of political messaging on social media, enabling researchers, policymakers, and the public to better understand emerging narratives, polarization dynamics, and the moral underpinnings of digital political communication.
Introduction
Modern political campaigns have been fundamentally transformed by the rise of social media platforms. Platforms such as Meta have become critical battlegrounds for political entities, enabling them to target specific voter segments with tailored messages. Chu et al. (2024) has shown that voters respond more favorably to ads on issues they personally prioritize, irrespective of partisanship. Understanding these targeting strategies is thus essential for both researchers and policymakers. As the volume and velocity of such dynamic content continues to grow, there is an urgent need for scalable tools that can systematically analyze these messages.
@@ Introduction
Manual annotation of large-scale text corpora is expensive and time-consuming, prompting widespread use of topic modeling techniques such as Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) and non-negative matrix factorization (NMF) (Lee and Seung 1999) for unsupervised topic discovery. Although effective in identifying topics, these methods produce topic representations that are not readily interpretable and often require human labeling. To automate this step, previous works have explored probabilistic labeling (Magatti et al. 2009), summarization-based methods (Wan and Wang 2016; Cano Basave, He, and Xu 2014), and the sequence-to-sequence model (Alokaili, Aletras, and Stevenson 2020). Recently, hierarchical topic modeling frameworks such as BERTopic (Grootendorst 2022) have gained wide adoption. BERTopic builds on contextualized representations and density-based clustering. Building on this approach, we propose a multi-step topic generation framework that leverages large language models (LLMs) (Brown et al. 2020) to produce semantically rich and coherent labels.
@@
Recent studies have also explored the use of LLMs to identify nuanced topics and uncover latent themes and arguments. LLMs-in-the-loop approach has been used for understanding fine-grained topics (Islam and Goldwasser 2025c, b). Islam and Goldwasser (2025b) employed a seed set of initial themes to guide their framework, while Islam and Goldwasser (2025c) assumed a predefined set of themes and then employ their framework to uncover arguments. In contrast, our method operates without any initial seed set, instead iteratively constructing a topic taxonomy from scratch. This enables greater flexibility and adaptability in capturing latent discourse structures within large, dynamic social media corpora.
@@
In this paper, we propose a novel two-pass iterative topic generation framework (Figure 1) that combines the strengths of unsupervised machine learning with the interpretive capabilities of LLMs. Our approach applies unsupervised clustering techniques to identify latent topical structures within a document corpus, followed by an iterative process where LLMs evaluate existing topics and generate new ones as needed. This method allows for the discovery of semantically rich and coherent topics without relying on predefined labels or human intervention. The output of this process is a set of coherent and interpretable document clusters, which can be used for downstream tasks such as supervised classification.
@@
We demonstrate the effectiveness of this approach through a comprehensive case study [^1] Code and dataset will be released upon publication. of Meta political advertisements, taken one month before the 2024 US Presidential Election. We uncover a diverse political issue taxonomy that captures the topical landscape of the electoral campaign. We further extend our analysis by examining the moral framings underlying these ads, under the lens of Moral Foundation Theory (Haidt and Graham 2007; Haidt and Joseph 2004). By analyzing how different political actors segment and appeal to voter demographics, our work provides valuable insights into contemporary campaign tactics while offering a methodological contribution that bridges traditional topic modeling with the interpretive capabilities of LLMs. Our contributions include:
@@
    A flexible framework for interpretable topic modeling that requires no human intervention.
@@
    A dynamic taxonomy of political issues and moral foundations, generated from unlabeled data with no seed set.
@@
    Systematic analysis of targeting strategies in political social media advertising.
@@
    Curate and release the 2024 US presidential election dataset to support future research.
Related Work
@@
Social media campaigns are a powerful force in shaping public discourse (Islam 2025b, a; Islam and Goldwasser 2024; Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021; Goldberg et al. 2021), influencing elections (Ribeiro et al. 2019; Silva et al. 2020; Fulgoni, Lipsman, and Davidsen 2016), and mobilizing voters (Aggarwal et al. 2023; Teresi and Michelson 2015; Hersh 2015a). Research on political microtargeting (Tappin et al. 2023; Zuiderveen Borgesius et al. 2018; Hersh 2015b) has been a key focus, with studies examining how targeted ads can influence voter preferences and behavior (Hirsch et al. 2024).
@@
Researchers have explored the role of targeted messaging in elections and the strategic dissemination of campaign content across platforms (Hackenburg and Margetts 2024; Islam, Roy, and Goldwasser 2023). These studies underscore the growing need for scalable methods to analyze social media content at scale and in context, particularly during major political events such as presidential campaigns.
@@
However, the sheer volume of social media data presents a significant challenge for researchers and practitioners. Traditional manual coding methods are often impractical for large datasets, prompting widespread use of topic modeling techniques such as LDA across various domains, including social media (Chen et al. 2019; Ozer, Kim, and Davulcu 2016) and political discourse (Lahoti, Garimella, and Gionis 2018; Mathaisel and Comm 2021). These methods have been instrumental in uncovering topics in text data, but they often struggle to produce coherent and interpretable topics (Chang et al. 2009; Mei, Shen, and Zhai 2007).
@@
To address these issues, modern embedding-based topic models like BERTopic (Grootendorst 2022) utilize contextualized representations and density-based clustering, resulting in more coherent and interpretable topics. Our framework builds on these advances by following the density-based clustering approach of BERTopic, but relying on LLM inference to generate an interpretable topic taxonomy.
@@
Prompt-based methods such as TopicGPT (Pham et al. 2023) have been shown to outperform traditional topic modeling techniques in terms of coherence and interpretability. These methods leverage LLMs to synthesize topic taxonomies and assign labels directly, allowing for more nuanced and context-sensitive analysis. Notably, Gilardi, Alizadeh, and Kubli (2023) showed that LLMs match or exceed the performance of expert human annotators in labeling tasks.
@@
More recently, LLMs-in-the-loop approaches have been used for understanding fine-grained topics (Islam and Goldwasser 2025c, b). In their earlier work, Islam and Goldwasser (2025b) guided their framework using a seed set of initial themes. Building on this, Islam and Goldwasser (2025c) assumed a predefined set of themes and focused on uncovering underlying arguments. In contrast, our method operates without any initial seed set.
@@
Another growing area of interest is the moral framing of political messages, often analyzed through the lens of Moral Foundations Theory (MFT) (Haidt and Graham 2007; Haidt and Joseph 2004). Earlier work relied on manually annotated datasets to train classifiers that detect moral appeals in text (Pacheco, Islam et al. 2022; Roy, Pacheco, and Goldwasser 2021). Recently, Islam and Goldwasser (2025a) has leveraged LLMs to generate moral labels and explanations through inference, allowing for more scalable and nuanced analysis. We extends this line of research by employing LLMs to synthesize arguments at the cluster level and assign corresponding moral labels, resulting in a more contextualized and semantically grounded interpretation of political discourse.
@@
Density-based Clustering
The input documents are first embedded into a high-dimensional space using a pre-trained sentence embedding model, such as Sentence-BERT (Reimers and Gurevych 2019). To avoid the curse of dimensionality, we use UMAP (McInnes, Healy, and Melville 2020) to reduce the dimensionality of the embeddings to a more manageable size. UMAP is a non-linear dimensionality reduction technique that preserves the local structure of the data, making it suitable for high-dimensional data such as text embeddings.
We then use HDBSCAN (McInnes, Healy, and Astels 2017), a hierarchical density-based clustering algorithm, to group similar text data points together and identify topical structures in the data. This is particularly useful for unstructured text data, where the distribution of data points may not follow a standard distribution and noise may be present. For each cluster, we extract the five items with the highest membership probabilities, which we refer to as the cluster representatives.
Iterative Topic Synthesis
To generate a domain-specific label taxonomy, we use an LLM to synthesize labels from the clusters. The initial label list is set empty, or optionally initialized with a small seed set to guide the generation process. This list will be iteratively expanded by the LLM. If the initial list is empty, the LLM is prompted to generate a label for the first cluster.
For every other cluster, we prompt the LLM whether any of the existing labels are accurate annotations for the cluster representatives. We use constrained decoding (Beurer-Kellner, Fischer, and Vechev 2024) to force the LLM to only answer with “yes” or “no”. If the LLM answers “no”, we prompt it to generate a new label for the cluster. We then add this label to the set of labels and repeat the process for all clusters. This iterative process continues until all clusters have been processed.
Cluster Representative Labeling
Supervised Classification
For classification, we found SetFit (Tunstall et al. 2022) to be particularly effective for text classification tasks with limited labeled data. SetFit uses a two-step process: first, it fine-tunes a pre-trained sentence transformer model on the labeled data points using a contrastive loss function. Second, it trains a classification head to map the fine-tuned embeddings to the label space. This model is then used to efficiently annotate the remaining unlabeled documents.
Case Study
@@
Care/Harm: : It suggests that someone other than the speaker is worthy of compassion or is experiencing harm, grounded in the values of kindness, tenderness, and care.
Fairness/Cheating: Emphasizes justice, personal rights, and independence; involves comparing with other groups. Advocates for equal opportunity and resists those who benefit without contributing (“Free Riders”).
Loyalty/Betrayal: Based on the values of loyalty to one’s country and willingness to sacrifice for the group. Activated by a sense of unity and collective responsibility—“one for all, and all for one”.
@@
Table 2: Six basic moral foundations (Haidt and Graham 2007; Haidt and Joseph 2004).
Dataset
Framework Application
We apply the proposed framework to the ad texts to identify a taxonomy of political issues discussed in the ads. We embedded the ad body texts using the all-MiniLM-L6-v2 model, and removed all ads with embedding vectors too similar to another ad (cosine similarity >0.95>0.95). We then clustered these embeddings with HDBSCAN using a minimum cluster size of $15$. This resulted in $4510$ of the ads being assigned to $72$ clusters. The prompts can be found in the Appendix A.
We used a locally hosted Llama-3.2-3B model (Grattafiori et al. 2024) to generate the cluster labels. Experiments were conducted on a local machine using an 11th Gen Intel Core i7-11390H CPU @ 3.40GHz, without GPU acceleration. We began the topic synthesis process without any seed set, which the model expanded to $14$ total topics. For annoation, the model was additionally given the option to assign a cluster the label ‘other’ if none of the topics suited its context. Of this list, the LLM never assigned the topic ‘border security’ to any of the clusters. See Table 1 for the final list of topics and the number of clusters assigned to each topic.
Moral Foundations

### v2

# Iterative Topic Taxonomy Induction with LLMs: A Case Study of Electoral Advertising

@@ Abstract
Social media platforms play a pivotal role in shaping political discourse, but analyzing their vast and rapidly evolving content remains a major challenge. We introduce an end-to-end framework for automatically inducing an interpretable topic taxonomy from unlabeled text corpora. By combining unsupervised clustering with prompt-based inference, our method leverages large language models (LLMs) to iteratively construct a taxonomy without requiring seed sets (predefined labels) or domain expertise. We validate the framework through a study of political advertising ahead of the 2024 U.S. presidential election. The induced taxonomy yields semantically rich topic labels and supports downstream analyses, including moral framing, in this setting. Results suggest that structured, iterative labeling yields more consistent and interpretable topic labels than existing approaches under human evaluation, and is practical for analyzing large-scale political advertising data.
Iterative Topic Taxonomy Induction with LLMs: A Case Study of Electoral Advertising
Alexander Brady1thanks: These authors contributed equally. 1Department of Computer Science, ETH Zürich, Zurich, Switzerland bradya@ethz.ch          Tunazzina Islam211footnotemark: 1 2Department of Computer Science, Purdue University, West Lafayette, IN 47907, USA islam32@purdue.edu
@@
1 Introduction
Modern political campaigns have been fundamentally transformed by the rise of social media platforms. Platforms such as Meta (previously known as Facebook) have become critical battlegrounds for political entities, enabling them to target specific voter segments with tailored messages. Chu et al. (2024) have shown that voters respond more favorably to ads on issues they personally prioritize, irrespective of partisanship. Understanding these targeting strategies is thus essential for both researchers and policymakers. As the volume and velocity of such dynamic content continues to grow, there is an urgent need for scalable tools that can systematically analyze such messages at scale and with minimal human supervision.
@@ 1 Introduction
Manual annotation of large-scale text corpora is expensive and time-consuming, prompting widespread use of topic modeling techniques such as Latent Dirichlet Allocation (LDA) Blei et al. (2003) and non-negative matrix factorization (NMF) Lee and Seung (1999) for unsupervised topic discovery. Although effective in identifying topics, these methods produce topic representations that are not readily interpretable and often require human labeling. To automate this step, previous works have explored probabilistic labeling Magatti et al. (2009), summarization-based methods Wan and Wang (2016); Cano Basave et al. (2014), and the sequence-to-sequence model Alokaili et al. (2020). Recently, hierarchical topic modeling frameworks such as BERTopic Grootendorst (2022) have gained wide adoption. BERTopic builds on contextualized representations and density-based clustering. Building on this approach, we propose a multi-step topic generation framework that leverages large language models (LLMs) Brown et al. (2020) to produce semantically rich and coherent labels.
@@
Recent studies have also explored the use of LLMs to identify nuanced topics and uncover latent themes and arguments. LLMs-in-the-loop approach has been used for understanding fine-grained topics Islam and Goldwasser (2025d, b). Islam and Goldwasser (2025b) employed a seed set of initial themes to guide their framework, while Islam and Goldwasser (2025d) assumed a predefined set of themes and then employ their framework to uncover arguments. In contrast, our method operates without any initial seed set, instead iteratively constructing a topic taxonomy from scratch. This enables greater flexibility and adaptability when inducing topic structure in large, dynamic social media corpora.
@@
In this paper, we propose a novel two-pass iterative topic generation framework (Figure 1) that combines the strengths of unsupervised machine learning with the interpretive capabilities of LLMs. Our approach applies unsupervised clustering techniques to identify latent topical structures within a document corpus, followed by an iterative process where LLMs evaluate existing topics and generate new ones as needed. This method allows for the discovery of semantically rich and coherent topics without relying on predefined labels or direct human intervention. The output of this process is a set of coherent and interpretable document clusters, which can be used for downstream tasks such as supervised classification.
@@
We demonstrate the effectiveness of this approach through a comprehensive case study of Meta political advertisements, taken one month before the 2024 US Presidential Election [^1] Code and dataset are publicly available here: https://github.com/alexander-brady/llm-topic-synthesis. We uncover a diverse political issue taxonomy that captures the topical landscape of the electoral campaign. We further extend our analysis by examining the moral framings underlying these ads, under the lens of Moral Foundation Theory Haidt and Graham (2007); Haidt and Joseph (2004). By analyzing how different political actors segment and appeal to voter demographics, the case study illustrates the types of substantive insights enabled by the proposed framework. Our contributions include:
@@
    A general, seed-free, LLM-guided framework for interpretable topic taxonomy induction from unlabeled corpora without human intervention.
@@
    An iterative topic synthesis mechanism that yields more consistent and better-aligned topic labels than classical topic models and single-shot LLM labeling under human evaluation.
@@
    A real world political advertising case study used to validate the framework and illustrate its analytical affordances.
@@
    Release of a curated dataset of 2024 U.S. presidential election ads to support future research.
2 Related Work
@@
Social media campaigns are a powerful force in shaping public discourse (Islam, 2025b, a; Islam and Goldwasser, 2025c; Islam et al., 2023a, b; Islam and Goldwasser, 2022; Capozzi et al., 2021; Goldberg et al., 2021), influencing elections (Ribeiro et al., 2019; Silva et al., 2020; Fulgoni et al., 2016), and mobilizing voters (Aggarwal et al., 2023; Teresi and Michelson, 2015; Hersh, 2015a). Research on political microtargeting Tappin et al. (2023); Zuiderveen Borgesius et al. (2018); Hersh (2015b) has been a key focus, with studies examining how targeted ads can influence voter preferences and behavior (Hirsch et al., 2024).
@@
Researchers have explored the role of targeted messaging in elections and the strategic dissemination of campaign content across platforms Hackenburg and Margetts (2024); Islam et al. (2023a). These studies underscore the growing need for scalable methods to analyze social media content at scale and in context, particularly during major political events such as presidential campaigns.
@@
However, the sheer volume of social media data presents a significant challenge for researchers and practitioners. Traditional manual coding methods are often impractical for large datasets, prompting widespread use of topic modeling techniques across various domains, including social media Chen et al. (2019); Ozer et al. (2016) and political discourse Lahoti et al. (2018); Mathaisel and Comm (2021). These methods have been instrumental in uncovering topics in text data, but they often struggle to produce coherent and interpretable topics Chang et al. (2009); Mei et al. (2007).
@@
To address these issues, modern embedding-based topic models like BERTopic Grootendorst (2022) utilize contextualized representations and density-based clustering, resulting in more coherent and interpretable topics. Our framework builds on these advances by following the density-based clustering approach of BERTopic, but relying on LLM inference to generate an interpretable topic taxonomy.
@@
Prompt-based methods such as TopicGPT Pham et al. (2023) have been shown to outperform traditional topic modeling techniques in terms of coherence and interpretability. These methods leverage LLMs to synthesize topic taxonomies and assign labels directly, allowing for more nuanced and context-sensitive analysis. Notably, Gilardi et al. (2023) showed that LLMs match or exceed the performance of expert human annotators in labeling tasks.
@@
More recently, LLMs-in-the-loop approaches have been used for understanding fine-grained topics Islam and Goldwasser (2025d, b). In their earlier work, Islam and Goldwasser (2025b) guided their framework using a seed set of initial themes. Building on this, Islam and Goldwasser (2025d) assumed a predefined set of themes and focused on uncovering underlying arguments. Lam et al. (2024) developed a concept induction algorithm using LLM with human-guided abstraction called LLooM, which has a seed operator for accepting user-provided seed term/set. In contrast, our method operates without any initial seed set.
@@
Another growing area of interest is the moral framing of political messages, often analyzed through the lens of Moral Foundations Theory (MFT) Haidt and Graham (2007); Haidt and Joseph (2004). Earlier work relied on manually annotated datasets to train classifiers that detect moral appeals in text Pacheco et al. (2022); Roy et al. (2021). Recently, Islam and Goldwasser (2025a) has leveraged LLMs to generate moral labels and explanations through inference, allowing for more scalable and nuanced analysis. We extend this line of research by employing LLMs to synthesize arguments at the cluster level and assign corresponding moral labels, resulting in a more contextualized and semantically grounded interpretation of political discourse.
@@ 3 Methodology
3.1 Density-based Clustering
The input documents are first embedded into a high-dimensional space using a pre-trained sentence embedding model, such as Sentence-BERT (Reimers and Gurevych, 2019). To avoid the curse of dimensionality, we use UMAP (McInnes et al., 2020) to reduce the dimensionality of the embeddings to a more manageable size. UMAP is a non-linear dimensionality reduction technique that preserves the local structure of the data, making it suitable for high-dimensional data such as text embeddings.
We then use HDBSCAN McInnes et al. (2017), a hierarchical density-based clustering algorithm, to group similar text data points together and identify topical structures in the data. This is particularly useful for unstructured text data, where the distribution of data points may not follow a standard distribution and noise may be present. For each cluster, we extract the five items with the highest membership probabilities, which we refer to as the cluster representatives.
Algorithm 1 Iterative Topic Taxonomy Induction
0: Document clusters $\mathcal{C}=\{C_{1},\dots,C_{n}\}$, cluster representatives $R(C_{i})$, LLM $\mathcal{M}$
0: Topic taxonomy $\mathcal{T}$ and cluster-topic assignments
1: Initialize topic list $\mathcal{T}\leftarrow\emptyset$
2:for each cluster $C_{i}\in\mathcal{C}$ do
3:  if $\mathcal{T}=\emptyset$ then
4:   Prompt $\mathcal{M}$ to generate a topic $t$ from $R(C_{i})$
5:   $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
6:  else
7:   Prompt $\mathcal{M}$ with $R(C_{i})$ and topics $\mathcal{T}$
8:   Constrain output to {yes, no}
9:   if no then
10:    Prompt $\mathcal{M}$ to generate a new topic $t$ from $R(C_{i})$
11:    $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
12:   end if
13:  end if
14:end for
15:for each cluster $C_{i}\in\mathcal{C}$ do
16:  Prompt $\mathcal{M}$ with $R(C_{i})$ and topics $\mathcal{T}$
17:  Constrain output to select exactly one topic $t\in\mathcal{T}$
18:  Assign topic $t$ to cluster $C_{i}$
19:end for
20:return $\mathcal{T}$ and cluster-topic assignments
3.2 Iterative Topic Synthesis
To generate a domain-specific label taxonomy, we use an LLM to synthesize labels from the clusters. The initial label set is empty. This list will be iteratively expanded by the LLM. If the initial list is empty, the LLM is prompted to generate a label for the first cluster.
For every other cluster, we prompt the LLM whether any of the existing labels are accurate annotations for the cluster representatives. We use constrained decoding Beurer-Kellner et al. (2024) to force the LLM to only answer with “yes" or “no". If the LLM answers “no", we prompt it to generate a new label for the cluster. We then add this label to the set of labels and repeat the process for all clusters. This iterative process continues until all clusters have been processed.
We formalize the iterative topic synthesis and annotation procedure in Algorithm 1.
3.3 Cluster Representative Labeling
3.4 Supervised Classification
For classification, we found SetFit (Tunstall et al., 2022) to be particularly effective for text classification tasks with limited labeled data. SetFit uses a two-step process: first, it fine-tunes a pre-trained sentence transformer model on the labeled data points using a contrastive loss function. Second, it trains a classification head to map the fine-tuned embeddings to the label space. This model is then used to efficiently annotate the remaining unlabeled documents.
@@ 3.4 Supervised Classification
4 Case Study
@@
4.1 Dataset
4.2 Framework Application
We apply the proposed framework to the ad texts to identify a taxonomy of political issues discussed in the ads. We embedded the ad body texts using the all-MiniLM-L6-v2 model, and removed all ads with embedding vectors too similar to another ad (cosine similarity >0.95>0.95). We then clustered these embeddings with HDBSCAN using a minimum cluster size of $15$. This resulted in $4510$ of the ads being assigned to $72$ clusters. The prompts can be found in the App.  A.
We used a locally hosted Llama-3.2-3B model Grattafiori et al. (2024) to generate the cluster labels. Experiments were conducted on a local machine using an 11th Gen Intel Core i7-11390H CPU @ 3.40GHz, without GPU acceleration. We began the topic synthesis process without any seed set, which the model expanded to $14$ total topics. For annotation, the model was additionally given the option to assign a cluster the label ‘other’ if none of the topics suited its context. Of this list, the LLM never assigned the topic ‘border security’ to any of the clusters. See Table 1 for the final list of topics and the number of clusters assigned to each topic.
Care/Harm: It suggests that someone other than the speaker is worthy of compassion or is experiencing harm, grounded in the values of kindness, tenderness, and care.
Fairness/Cheating: Emphasizes justice, personal rights, and independence; involves comparing with other groups. Advocates for equal opportunity and resists those who benefit without contributing (“Free Riders").
Loyalty/Betrayal: Based on the values of loyalty to one’s country and willingness to sacrifice for the group. Activated by a sense of unity and collective responsibility—“one for all, and all for one".
@@ 4.2 Framework Application
Table 2: Six basic moral foundations Haidt and Graham (2007); Haidt and Joseph (2004).
4.3 Moral Foundations